### v2

# Exclusion of a direct progenitor detection for the Type Ic SN 2017ein based on late-time observations

Abstract
To date, SN 2017ein is the only Type Ic supernova with a directly identified progenitor candidate. This candidate points to a very massive ($>$45 $M_{\odot}$) Wolf-Rayet progenitor, but its disappearance after the explosion of SN 2017ein remains unconfirmed. In this work, we revisit SN 2017ein in late-time images acquired by the Hubble Space Telescope (HST) at 2.4–3.8 yrs after peak brightness. We find this source has not disappeared and its brightness and color remain almost the same as in the pre-explosion images. Thus, we conclude that the pre-explosion source is not the genuine progenitor of SN 2017ein. It is not much likely to be a companion star of the progenitor, since it has a much lower extinction than SN 2017ein; its color also seems inconsistent with a star cluster, indicated by the newly added magnitude limit in F336W, apart from F [^5] W and F814W. We suggest, therefore, this source is an unrelated star in chance alignment with SN 2017ein. Based on the low ejecta mass, we propose that SN 2017ein is most likely originated from a moderately massive star with $M_{\rm ini}$ $\sim$ 8–20 $M_{\odot}$, stripped by binary interaction, rather than a very massive Wolf-Rayet progenitor.
1 Introduction
Core-collapse supernovae (CCSNe) are violent explosive phenomena associated with massive stars at the end of their lives. Among them, Type Ib SNe refer to those without H lines observed in their spectra, and Type Ic SNe are those without H and He lines. It is believed that their progenitors have the outer envelopes stripped; some researchers also propose that Type Ic SNe may still retain some He envelope, but the He lines are not apparent in the spectra due to a lack of excitation (Dessart et al. 2012). In the local Universe, Type Ib and Type Ic SNe constitute $\sim$20of all CCSNe (Smith et al. 2011a).
Theoretically, Type Ib/Ic SNe may arise from two possible progenitor channels (Woosley & Eastman 1997). In the single progenitor channel, a massive star with initial mass exceeding 25–30 $M_{\odot}$ can be stripped by its strong stellar winds (Conti 1978) and become a Wolf-Rayet (WR) star, characterized by extreme luminosity and prominent emission lines, before explosion. In the binary progenitor scenario, a star may be stripped by a close companion via stable or unstable mass transfer involving Roche-lobe overflow and common-envelope evolution (Podsiadlowski 2014). In the latter channel, the progenitors may be of lower initial masses compared with that of a single WR progenitor (Podsiadlowski et al. 1992).
There have been direct progenitor detections in the pre-explosion images for $\sim$20 Type II-P SNe (e.g.
SN 2005cs, Maund et al. 2005; SN 2017eaw, Rui et al. 2019; Van Dyk et al. 2019; SN 2023ixf, Niu et al. 2023), 6 Type IIn SNe (e.g. SN 2005gl, Gal-Yam et al. 2007; SN 2010jl, Smith et al. 2011b; Fox et al. 2017; Niu et al. 2024a) and 6 Type IIb SNe (e.g. SN 2008ax, Crockett et al. 2008; Folatelli et al. 2015; SN 2017gkk, Niu et al. 2024b). In contrast, the search for Type Ib/Ic SN progenitors proves to be extremely difficult. Since their outer envelopes are stripped, the progenitors’ spectral energy distributions (SEDs) peak at ultraviolet instead of optical wavelengths (Maund 2018). In addition, Type Ib/Ic SNe are often found in young star-forming regions with very high interstellar extinctions (Sun et al. 2023). Up to now, only two progenitors of Type Ib SNe have been directly detected (i.e. iPTF13bvn, e.g. Cao et al. 2013; Bersten et al. 2014; SN 2019yvr, Kilpatrick et al. 2023; Sun et al. 2022), and only one progenitor candidate detection has been reported for a Type Ic SN (i.e. SN 2017ein, which shall be discussed later in this work; Kilpatrick et al. 2018 (K18), Van Dyk et al. 2018 (VD18), and Xiang et al. 2019 (X19)).
The difficulty in direct progenitor detections motivates researchers to use alternative methods to determine the dominant evolutionary channel for the Type Ib/Ic SN progenitors. For example, Smith et al. (2011a) argued that the occurrence rate of Type Ib/Ic SNe is too high to be compatible with the small number of massive WR stars at the high-mass end of the stellar initial mass function. Lyman et al. (2016), Taddia et al. (2018) and Woosley et al. (2021) analyzed the light curves of stripped-envelope SNe (SESNe; Types IIb, Ib and Ic), all supporting interacting binaries as the dominant progenitor channel for Type Ib/Ic SNe. By analyzing the nebular-phase spectra and SN environments, Fang et al. (2018) and Sun et al. (2023) suggested that Type Ic SN progenitors may undergo a hybrid envelope-stripping mechanism, in which the hydrogen envelopes are stripped by binary interaction while the helium envelopes are stripped by intense stellar winds. Solar et al. (2024) confirmed observationally for a statistically large sample that Type Ic SN progenitors are binary systems, using molecular clouds in the environments of the SN explosions as tracer of lifetimes.
SN 2017ein is the only Type Ic supernova with a progenitor detection. First discovered by Ron Arbour, it exploded in May 2017 in the nearby spiral galaxy NGC 3938, with a distance of $\sim$ 17 Mpc (Tully et al. 2009). K18, VD18 and X19 conducted analyses of SN 2017ein. They found its spectrum lacked hydrogen and helium lines, consistent with a Type Ic classification; the peak magnitude in the V band is $-16\sim-17.5$ mag, dimmer than 80% of Type Ic SNe. These characteristics make it very similar to the carbon-rich, low-luminosity Type Ic SN 2007gr (Chen et al. 2014). Although the Galactic extinction of SN 2017ein is quite small of only $A_{V}$ = 0.058 mag (Schlafly & Finkbeiner 2011), these studies found its host-galaxy extinction to be much larger, with estimated values ranging from $A_{V}$ = 0.94 to 1.84 mag, as measured with a variety of methods.
In the pre-explosion images of the Hubble Space Telescope (HST) , K18, VD18, and X19 found a point source located exactly at the SN position. They proposed that it is most likely to be the progenitor of SN 2017ein. VD18 also suggested that it may be the binary companion or host star cluster of this SN. Any of these scenarios would suggest a very massive WR progenitor of > 45 $M_{\odot}$. The ejecta mass of SN 2017ein was, however, only $\sim$ 1 $M_{\odot}$ (VD18 and X19), which seems too low to be consistent with such a high progenitor mass (Lyman et al. 2016, Taddia et al. 2018). In addition, Jung et al. (2022) calculated the synthetic SEDs of massive WR progenitors with strong winds; they found the predicted colors are significantly different from that of the reported progenitor candidate of SN 2017ein.
It is very important to confirm whether the reported candidate was the genuine progenitor by inspecting whether it has disappeared at late times (Maund et al., 2015; Niu et al., 2024a). In this work, we carry out such a study for the progenitor candidate of SN 2017ein and re-examine its object type and relation with the SN. This paper is organized as follows. Section 2 describes the data used in this work and how we conducted photometry. In Section 3, we present and discuss the results. This paper is then closed with a summary in Section 4.
2 Data and photometry
The data used in this work are from observations conducted with the HST (see Table 1 for a complete list). The pre-explosion images were observed with the Wide Field Planetary Camera 2 (WFPC2) Planetary Camera (PC), while the late-time images were acquired with the Wild Field Camera 3 (WFC3) Ultraviolet-Visible (UVIS) channel at t = 2.4–3.8 years after the peak brightness. All the observations were conducted in the F [^5] W and F814W filters, with an addition of F336W at t = 2.5 yrs. Besides, an F438W image, taken right after the SN explosion, was used as a reference image for differential astronomy. The HST data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/7n8j-cy13 (catalog https://doi.org/10.17909/7n8j-cy13).
We retrieved the data from the Barbara A. Mikulski Archive for Space Telescopes (MAST; https://mast.stsci.edu). For the pre-explosion WFPC2 images, we used the iraf task crrej (Gonzaga 2002) to remove cosmic rays. For the late-time WFC3 images, we used the tweakreg (Avila et al. 2015) package to align the individual exposures (*_flc.fits images) and combined them with the astrodizzle (Hack et al. 2012) package to remove the cosmic rays.
The dolphot (Dolphin 2000) package was used to conduct point-source photometry. In order to see how the photometry parameters may influence the final results, we tested the use of different parameters of FitSky = 2 or 3, RAper = 3 or 8, ApCor = 0 or 1, Force1 = 0 or 1. We find that the results do not change significantly with the chosen values. In this work, we adopted the result obtained with FitSky = 2, RAper = 3, ApCor = 1, Force1 = 1, as recommended for crowded fields. We also used common stars to correct for the possible systematic errors between observations at different epochs; the details are provided in Appendix A.
Table 1: HST data and photometry.
Date (UT) Time from Instrument Filter Exposure Program Systematic error in Magnitudec
peak (yr) Time (s) ID photometryb (mag) (mag)
2007 Dec 12.4 $-10$ WFPC2/PC F [^5] W 460 10877 --0.37 (0.01) 24.75 (0.11)
2007 Dec 12.4 $-10$ WFPC2/PC F814W 700 10877 --0.26 (0.01) 24.56 (0.18)
2017 Jun 12.1 00 WFC3/UVIS F438W 270 14645 d
2019 Nov 7.7 $2.4$ WFC3/UVIS F [^5] W 1913a 15691 --0.03 (0.004) 24.83 (0.05)
2019 Nov 7.8 $2.4$ WFC3/UVIS F814W 1272a 15691 --0.08 (0.006) 24.58 (0.14)
2019 Dec 1.0 $2.5$ WFC3/UVIS F336W 2625 15853 23.61 (0.06)
2019 Dec 7.1 $2.5$ WFC3/UVIS F [^5] W 1913a 15691 --0.05 (0.004) 24.77 (0.06)
2019 Dec 7.2 $2.5$ WFC3/UVIS F814W 2544 15691 --0.06 (0.005) 24.50 (0.05)
2021 Apr 2.9 $3.8$ WFC3/UVIS F [^5] W 710 16239 24.72 (0.07)
2021 Apr 2.9 $3.8$ WFC3/UVIS F814W 780 16239 24.76 (0.14)
$a$ Some exposures are discarded due to bad image quality and are not included in the exposure time as listed here.
$b$ The values are relative to observations on $t=3.8$ yrs.
$c$ The systematic errors have been corrected except for F336W.
$d$ This image is only used as a reference image for differential astrometry.
3 Results and discussion
Figure 1 shows the site of SN 2017ein in the pre-explosion and late-time observations. For better determination of the SN position, differential astrometry was used to align the images with respect to the image taken near peak brightness. We selected more than ten common stars as reference stars, and the uncertainty of the alignment reaches 0.7 pixel. It is clearly shown that in the pre-explosion images, there is a source at the SN position in both F [^5] W and F814W bands, with a positional difference of only $\sim$0.13 pixel. This source corresponds to the progenitor candidate of SN 2017ein as reported in the previous studies (K18, VD18, and X19). For convenience, we shall refer to it as Source A hereafter. In the late-time images, however, Source A does not disappear and the SN position remains still bright. The sharpness parameter of Source A reported by dolphot is very close to zero, which implies that Source A is a point source.
Figure 1:  Pre- and post-explosion HST images of the site of SN 2017ein (F336W is not included here). The SN position is denoted by the white crosshair. Each image has a size of $\sim$ 160$\times$160 pc and is oriented with North up and East to the left.
[…]
According to these phenomena, the most likely scenario is that the pre-explosion object, a star or star cluster, has not disappeared after the explosion of SN 2017ein since the late-time brightness and color are very consistent with the pre-explosion ones. The extra light in F814W and the red color at $t$ = 2.5 yrs could arise from a light echo (Boffi et al. 1999; Bond et al. 2003) or CSM interaction (Pellegrino et al. 2022; Fraser 2020), which has completely faded by the last epoch. A light echo or CSM interaction could also be present at $t$ = 2.4 yrs, but the observations have such large photometric uncertainties that we could not identify their contribution. It is noteworthy that another scenario is also considered, where the SN ejecta shocks the surviving binary companion, causing it to be over luminous and inflated for several years after the SN explosion (Hirai et al., 2018; Ogata et al., 2021; Chen et al., 2023). However, in such a scenario, there is little possibility that Source A’s magnitudes of different bands remain this consistent with the pre-explosion ones.
Figure 2:  Brightness and color evolution of Source A at different epochs. The pre-explosion magnitudes measured by K18, X19 and VD18 are also plotted for comparison. The dashed lines represent the inverse-variance weighted averages of the pre-explosion and the last epoch, and the green-shaded bands indicate the errors. All the magnitudes have been corrected for systematic bias between the different epochs.
We analyze Source A on the (F336W--F [^5] W, F [^5] W--F814W) color-color diagram (Figure 3) to determine whether it is a star or a star cluster. To reduce the photometric uncertainties in F [^5] W/F814W, we use the inverse-variance weighted averages of the pre-explosion magnitude and that at the last epoch as the adopted magnitude of Source A (i.e. 24.73 $\pm$ 0.06 mag and 24.69 $\pm$ 0.11 mag, respectively). For F336W, we conservatively regard the derived magnitude of 23.61 $\pm$ 0.06 mag as a brightness upper limit for the underlying star or star cluster, since this magnitude is measured at t = 2.5 yrs, by which time there could still be some extra light caused by a light echo or CSM interaction. This possible extra light may also influence the brightness in F [^5] W and F814W, but the effects are believed to be negligible (see above discussion). For comparison, we show synthetic colors for stars of different effective temperatures (based on the Castelli & Kurucz 2004 stellar atmospheric models) and star clusters of different ages (based on the bpass model population synthesis; Eldridge et al. 2017; Stanway & Eldridge 2018; Byrne et al. 2022). For both stellar and cluster SEDs, we also apply a range of different extinctions.
In Fig. 3(a), the color of Source A appears very different from that expected for star clusters. The previous studies (K18, VD18, and X19) also discussed the possibility of Source A being a young star cluster but only with the F [^5] W -- F814W color from the pre-explosion observations. Now with the addition of the F336W limit, we suggest this scenario is not very likely since the young star clusters would be too bright in this filter. This analysis, however, is based on a standard extinction law of $R_{V}$ = 3.1; the color of Source A could still be consistent with that of a star cluster in case of a different host-galaxy extinction law. Moreover, since Source A lies quite close to the cluster models on the diagram, we might also consider the possibility of it truly being a cluster, with an age of $10^{6.5}$ yrs and a total extinction of $A_{V}=0.3$ mag. We estimated the mass of such a cluster to be $10^{3}$ $M_{\odot}$. Binaries or single stars with $M<114M_{\odot}$ can explode at this age, based on BPASS models. It is noteworthy that this extinction value is much lower than that of SN 2017ein (1.26 $\pm$ 0.2 mag, K18; 1.0–1.9 mag, VD18; 1.29 $\pm$ 0.18 mag, X19; measured with different methods), which might be explained by the supposition that the SN explodes from a dusty star within the cluster. Moreover, we find the influence of metallicity on the cluster models is negligible in our work.
[…]
Figure 3:  F336W--F [^5] W v.s. F [^5] W--F814W color-color diagrams for Source A (red circle) and model stars (bottom panel, solid lines) and star clusters (top panel, solid lines). The arrows in the upper-left corner show the reddening vector under a standard extinction law with $R_{V}$ = 3.1.
Jung et al. (2022) noticed that Source A is too blue and too bright to align with any of the stellar atmospheric models that they developed for single WR progenitors of Type Ic SNe. This can now be explained by our work, as it turns out that Source A is not the genuine progenitor but most likely an unrelated star in chance alignment.
K18, VD18, and X19 suggested that Source A could be the progenitor of SN 2017ein, or its companion star, or a host star cluster. In any scenario, Source A would support a very massive WR progenitor with an initial mass of $>$45 $M_{\odot}$. In this work, however, we find Source A is actually an unrelated star in chance alignment. Based on the low ejecta mass ($\sim$1 $M_{\odot}$; VD18; X19), we suggest the progenitor of SN 2017ein is most likely to be a moderately massive star of 8–20 $M_{\odot}$ stripped in a binary system.
4 Summary
In this work, we re-investigate the progenitor candidate reported for SN 2017ein with the new HST images acquired at 2.4–3.8 yrs after peak brightness. We find that the candidate has not disappeared, with its brightness and color almost the same as the pre-explosion ones. Thus, we conclude that the candidate is not the genuine progenitor. It is neither a companion star of the progenitor considering its much lower extinction than SN 2017ein. The color-color diagram analysis also indicates that this source is less likely to be a star cluster. Therefore, we suggest it is most likely to be an unrelated star in chance alignment with SN 2017ein. Based on the low ejecta mass, we suggest that the progenitor is more likely to be a moderately massive star stripped in a binary, rather than a very massive WR star as suggested by previous works.
acknowledgments
This work is supported by the Strategic Priority Research Program of the Chinese Academy of Sciences, Grant No. XDB0550300. NCS’s research is funded by the NSFC grants No. 12303051 and No. 12261141690 and ZXN acknowledges support from the NSFC through grant No. 12303039. JFL acknowledges support from the NSFC through grants No. 11988101 and No. 11933004 and from the New Cornerstone Science Foundation through the New Cornerstone Investigator Program and the XPLORER PRIZE. We thank the anonymous referee for the helpful comments.
Appendix A Systematic error correction
[…]
Figure A1:  Comparison of photometry as reported by dolphot for observations at different epochs. The black dashed line shows the one-to-one relation while the orange dash-dotted line corresponds to systematic shifts as calculated by inverse-variance weighted averages.
Figure A2:  Histograms of magnitude differences with respect to the epoch of t = 3.8 yrs. The solid lines are Gaussian fit curves and the dashed lines represent the means of the Gaussian fit. Normalization has been applied to each graph.